OpenStudy (anonymous):

simplify the following radicals please show work 1. 3√14 * -2√21 2. √2(√8 + √6)

OpenStudy (lieutenantgeneral):

3√14 * -2√21 = -46√6 √2(√8 + √6) = √2 (2 √2 + √6)

OpenStudy (anonymous):

could you show your work

OpenStudy (lieutenantgeneral):

x^2 - 8x + 4 is the mini polynomial for √2(√8 + √6)

OpenStudy (lieutenantgeneral):

2 (2 + √3) 4 + 2 √3

OpenStudy (anonymous):

is that for the first one??

OpenStudy (anonymous):

1. 3√14 * -2√21 3√14=3*√7*√2 -2√21=-2*√7*√3 =3*√7*√2*-2*√7*√3=7*3√3*-2√2=42√6 2.√2(√8 + √6) √8=2√2 √6=√2√3 =√2(2√2+√2√3) =√2√2(2+√3)=2(2+√3)